OpenStudy (anonymous):

Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3 Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2. Then use the _________________________ to subtract 1 from bot

OpenStudy (anonymous):

Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3 Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2. Then use the _________________________ to subtract 1 from both sides of the equation. This gives the solution x = –3. Therefore, given the equation 3x + 1 = 2 + 2x – 4, x is equal to –3. Which justification was left out of the paragraph proof above?

OpenStudy (anonymous):

You are subtracting by 1 both the sides equally: So here the property will be Subtraction Property of equality..

OpenStudy (anonymous):

are the best medal 4 u!!!!

OpenStudy (anonymous):

you are the best